The marginal cost function of a firm is 2 + 3e^x Find the



total and average cost functions if the fixed cost is Rs. 500.

1
Expert's answer
2022-03-25T15:07:32-0400

Total cost is TC=FC+MC×Q=500+(2+3e^x)×Q;

Average cost is AC=TC/Q=500/Q+2+3e^x